Hi!

Can someone please help me with the following question? If you can show all your steps so I can understand the solution I would really appreciate it!
Thanks :)

2- Below you can find scores of a test: {56,67,67,70,73,73,75,76,77,79,80,81,83,83,85,87,88,89,92,93 } (a) Obtain a frequency​​​​​​​
Answer #1

2.

(a)

We take classes as 50-59, 60-69, 70-79, 80-89, 90-99 as so frequency distribution is as follows.

Class limits Class boundaries Class mark (xi) Frequency (fi)
56-60 55.5-60.5 58 1
61-65 60.5-65.5 63 0
66-70 65.5-70.5 68 3
71-75 70.5-75.5 73 3
76-80 75.5-80.5 78 4
81-85 80.5-85.5 83 4
86-90 85.5-90.5 88 3
91-95 90.5-95.5 93 2
Total - 20

(b)

The frequency histogram is as follows.

Histogram of x 4 3 Frequency 2 60 70 80 90

(c)

Mean is given by

- Σ= fiti - 1575 = 78,75

Median is the 20/2 = 10th element in increasing order. So, it is the 3rd element of the class 76-80 (median class) and so median can be obtain as follows.

T =1+ 1/2-C *h

where

l = lower limit of median class interval = 76

C= cumulative frequency of preceding to the median class frequency = 1+0+3+3 = 7

f = frequency of the class interval to which median belongs = 4

h = width of the class interval = 80.5-75.5 = 5

N= total frequency = 20

...0 = 76 + 10 – 7 -*5= 79.75

Variance is given by

2 Σ= f(x; – 1)? _ 1825 - 91.25 20

(d)

Box plot is as follows.

160 7 0 80 90

(e)

For 56, standard score is given by

56 – 78.75 -= -2.381579 91.25

For 92, standard score is given by

92 – 78.75 -= 1.387073 91.25

(f)

Mean = 70 and standard deviation = 10

For 56, standard score is given by

56 – 70 -1.4 10

For 70, standard score is given by

70-70 - = 0 10

For 92, standard score is given by

92 – 70 102.2